# revtex4-2Repair the float

### v1

@@ Abstract
We study the physics of the strong-coupling Hubbard model in a kagome lattice ribbon under mechanical tension and half-filling. It is known that in the absence of strain, the lattice symmetry of the system and strong electronic interactions induce magnetic frustration. As uniaxial strain is applied, the ribbon exhibits various configurations with energy oscillations that depend on the direction of the strain axis. The ground states are obtained by density matrix renormalization group calculations. We find that the system is characterized by strong antiferromagnetic bonds distributed throughout the lattice along directions and patterns that depend on the strain directions and may coexist with easily polarizable sites that are only weakly correlated to their neighbors. We identify frustration and correlation measures that follow the strain and interaction dependence of the system well. These results illustrate that strain-dependent magnetic susceptibility could be explored experimentally to help probe the role of symmetry and interactions in these systems.
@@ I Introduction
The quantum spin liquid is an exotic and fascinating phase of matter [1, 2, 3]. Proposed by P. W. Anderson in 1973 [4], he suggested that such phase could emerge as the ground state of a spin-1/2 antiferromagnet on geometrically frustrated triangular lattices [5]. However, after years of debate on whether a quantum spin-liquid (QSL) phase would emerge in a triangular antiferromagnet, detailed studies suggest this is not the case (see [1] and references therein). Indeed, the search for the realization of this elusive phase of matter has elicited great effort of the scientific community over the years [6, 7, 8, 9, 10].
@@
Different phases have been predicted theoretically in frustrated kagome lattices using antiferromagnetic Heisenberg and/or Hubbard models [41]. While the former accounts strictly for the spin degrees of freedom in the system, the latter also accounts for charge fluctuations and can then describe a wider range of phenomena. Schnyder et al. [17] showed that different ground states could be obtained by controlling the spatial anisotropy of antiferromagnetic Heisenberg couplings on a kagome lattice, resulting in spiral phases and possible noncoplanar order. More recently, Nayga and Vojta [42] have also shown that the ground state of a kagome magnet with classical spins can be controlled by applied mechanical distortion of the lattice, including the finding that triaxial strain results in a noncoplanar spin-liquid phase. These studies point out that the magnetic properties of kagome systems are very sensitive to changes in the geometric frustration of the lattice. The reduction in symmetry would then be expected to strongly affect systems that incorporate charge fluctuations, such as those present in kagome metals. The Hubbard model provides a natural description, as the relative hoppings can be seen as effective control parameters that allow one to address the important question of how strain modifies the ground state properties. Sun and Zhu [43] have shown recently that by controlling the Coulomb repulsion in a Hubbard model of a kagome ribbon it is possible to drive the system across multiple quantum phase transitions. This includes a strongly interacting regime where the ground state is effectively governed by an antiferromagnetic Heisenberg model and displays QSL characteristics. For weaker interaction, these authors see evidence for different correlated phases in this geometrically frustrated system. We are interested in exploring the role that strain plays in the behavior of kagome systems as interactions change their response from the metallic to the insulating regime. The question we want to explore is how strain modifies possible frustrated ground states in the different regimes with varying correlations and charge fluctuations.
@@
In this work, we investigate the effect of a uniaxial uniform strain in a kagome nanoribbon. An schematic representation of the system is depicted in Fig. 1. Adopting a single-band Hubbard model, the effective role of strain is to modify the strength of the hopping matrix elements between electronic orbitals localized in the lattice sites. This generally results in anisotropic couplings along the three lattice vectors, thus reducing geometric frustration in the system. By performing density matrix renormalization group (DMRG) calculations within a tensor networks platform [44, 45], we compute the ground state properties of the system, including the spatial distribution of spin-spin correlations and the local magnetization upon weak magnetic fields applied. Our results show that the anisotropy produced by the applied strain induces the formation of Neel lines of strongly antiferromagnetically correlated bonds arranged on different patterns and along different lattice directions, depending on the strain orientation. This behavior can be understood as arising from the suppression of the geometrical frustration in the system. Interestingly, we also find that scattered among the well-defined Neel lines with antiferromagnetic correlations, sites exist that are only weakly coupled to their neighbors. These ‘interstitial’ sites are therefore easily polarizable by weak external magnetic fields, while the Neel lines remain in their antiferromagnetic coupled structure. The pattern of Neel bonds and spatial distribution of polarizable sites depend critically on strain orientation and becomes more pronounced for larger values of Coulomb interaction as the system approaches the kagome antiferromagnet regime. To better characterize the response to strain, we introduce two intuitive quantities: a geometric indicator to quantify the structural frustration in the system and another that measures how local spin correlations reflect the degree of frustration in the ground state. These measures are good predictors of the different spin correlation profiles seen in the system as the strain and interaction change.
@@ II Model and numerical Methods
Figure 1: (a) Schematic representation of the kagome nanoribbon. Circles represent the atomic sites of a tight-binding model while solid lines correspond to nearest-neighbor hopping matrix elements. Dashed lines at the top indicate periodic boundaries, in which case the ribbon becomes a cylinder. $\hat{\bf e}_{1}$ and $\hat{\bf e}_{2}$ span the unit cell, while ${\boldsymbol{a}}_{j}$ ($j=1,2,3$) represent the nearest-neighbor vectors. The angle $\theta$ defines the direction of applied strain with respect to $\hat{\bf e}_{1}$. This figure corresponds to $L_{x}=4$ unit cells along the $x$ direction and $L_{y}=3$ unit cells along $\hat{\bf e}_{2}$. The dashed hexagon (magenta) contains a minimal unit cell and associated bonds used in the text.
@@
[…]
@@ III Numerical Results
We set $t=1$ as the energy unit, $\hbar=1$ hereafter, and ensure that the system is in the half-filling regime by setting $\varepsilon_{0}=-U/2$. We explore values of $U$ from $U=5$ to $U=20$, corresponding to different phases of the system. For large values of $U$, the system is well described by an antiferromagnetic Heisenberg model [43]. The number of sites in the kagome nanoribbon increases rapidly with $L_{x}$ and $L_{y}$, $N=L_{y}(3L_{x}+2)$; we focus our analysis on a system with $L_{x}=12$ and $L_{y}=3$, that under periodic boundary conditions corresponds to the YC6 cylinder in Ref. [25]. We have allowed DMRG link dimension up to 5000, resulting in truncation errors typically less than 10-6.
@@ III.1 Unstrained vs strained regimes
We first look at the spin-spin correlations in the system for a strain-free lattice, $\epsilon=0$. As defined above, we use ${\cal C}_{\rm link}$ to provide a convenient visualization of the correlations. Figure 2 shows a heat map of ${\cal C}_{\rm link}$ for different Hubbard interaction values, placing the system into different phases, as follows: $U=5$ (a), metallic phase; $U=10$ (b), quantum spin liquid phase; and $U=20$ (c), kagome antiferromagnetic phase [43]. As the color bar shows, darker links correspond to stronger antiferromagnetic pairs. For small $U$, the correlation links are not as strong and are homogeneously distributed, as the weak interactions in this metallic regime would suggest. Despite the homogeneous distribution of hopping terms in the Hamiltonian, we notice some differences, likely due to edge effects on the ribbon. For larger $U$, the antiferromagnetic correlations appear stronger (darker links) and not as homogeneously distributed as the system approaches the Heisenberg antiferromagnetic regime. In an infinite system, the spin-dominated ground state is expected to have high frustration without strain. A spin liquid phase may even emerge, with not yet fully understood characteristics [25, 43]. We aim to study the role that strain may play in modifying the frustration otherwise present in the system. We will contrast Fig. 2 with the link correlation in the presence of strain to expose some of its effects on the ground state.
@@
Figure 2: Link correlations, ${\cal C}_{\rm link}$, in the absence of strain ($\epsilon=0$) for (a) $U=5$, (b) $U=10$ and (c) $U=20$. Darker links represent stronger antiferromagnetic alignments, clearly enhanced in panel (c). The uniform link distribution in (a) becomes more inhomogeneous for increasing $U$. Apart from strong links along the edges of this open ribbon, there appears to be no overall order of the antiferromagnetic bonds in (c).
@@
Figure 3: (a) Ground state energy per site, $E_{0}/N$, as function of strain direction $\theta$ for $\epsilon=0.1$ (a), $\epsilon=0.15$ (b), $\epsilon=0.2$ (c) and $\epsilon=0.3$ (d). Blue and red curves correspond to open and periodic boundary conditions, respectively. The black dashed line shows $\zeta(\theta)$ defined in Eq. (5) and scaled as per [47]. In all panels $U=5$, i.e., in the metallic regime for the unstrained system.
@@
We now analyze the effect of strain on the link-correlation pattern over the ribbon. Because the effect of strain on the correlations depends on the angle $\theta$ along which the system is stretched, it is instructive to analyze first how the ground state energy $E_{0}$ changes with $\theta$. Figure 3 shows $E_{0}(\theta)$ vs $\theta$ for $U=5$ and increasing strain values. Blue circles and red squares correspond to open (OBC) and closed boundary conditions (PBC), respectively. In all cases, $E_{0}$ exhibits an oscillatory behavior with clear maxima and minima as $\theta$ varies from 00 to $\pi$, as well as an overall increasing value for larger strains. The oscillations with strain direction are better defined for $\epsilon=0.30$ [Fig. 3(d)] but already visible for $\epsilon=0.15$ [see Fig. 3(b)]. The minima occur for $\theta\simeq\pi/6$, $\pi/2$ and $65\pi/$ 6, which correspond to strain applied orthogonal to one of the lattice vectors ${\boldsymbol{a}}_{j}$. In contrast, the $E_{0}$ maxima are observed around $\theta=0$, $\pi/3$ and $32\pi/$ 3, for which the strain direction lies along one of the lattice vectors. The general behavior of $E_{0}$ increasing with $\epsilon$ can be seen to be associated with the suppression of antiferromagnetic correlations, which contribute to lowering the energy of the system. This is consistent with the fact that under PBC $E_{0}$ is lower, as additional singlet-like correlations appear across the edges. We also notice that $E_{0}(\theta)$ is not fully symmetric around $\theta=\pi/2$, a direct consequence of the slanted geometry that defines the ends of the ribbon. Such asymmetry is even more pronounced under PBC, as seen in the red curves of Fig. 3.
@@
This quantity can be scaled and compared with the ground state energy $E_{0}(\theta)$  [^1] The scaling factor is obtained as follows: with $\zeta_{max}\equiv\text{max}\left\{\zeta(\theta)\right\}$ and $\zeta_{min}\equiv\text{min}\left\{\zeta(\theta)\right\}$; the equivalent minimum (maximum) value of $E_{0}(\theta)$ for those angles in the OBC curve are $E_{min}$ ($E_{max}$). Then $\zeta\rightarrow\left(\frac{\zeta-\zeta_{min}}{\zeta_{max}-\zeta_{min}}\right)\left(E_{max}-E_{min}\right)+E_{min}$ gives the desired scale for $\zeta(\theta)$ shown in Fig. 3.. Figure 3 shows $\zeta(\theta)$ as a dashed black curve; evidently, the scaled function $\zeta(\theta)$ follows $E_{0}(\theta)$ quite well. As mentioned above, the results for PBC are more affected by end effects, and the agreement is poorer. Nevertheless, the positions of maxima and minima predicted by $\zeta(\theta)$ agree quite nicely with $E_{0}(\theta)$ for both OBC/PBC configurations obtained via DMRG. These results suggest that the maxima of $E_{0}(\theta)$ are obtained whenever the geometric frustration in the system is higher, even for the case $U=5$ (metallic regime when unstrained) shown in Fig. 3.
Figure 4: Link correlations for $\epsilon=0.30$ and different strain orientations: $\theta=0$ (a), $\theta=\pi/6$ (b), $\theta=\pi/3$ (c), $\theta=\pi/2$ (d), $\theta=2\pi/3$ (e) and $\theta=5\pi/6$ (f). The left (right) column panels correspond to maxima (minima) in the OPC $E_{0}(\theta)$ of Fig. 3. For energy maxima (left column), links with antiferromagnetic correlations form 2D oblique lattices. For energy minima (right column), links form 1D-like arrays with stronger (darker) antiferromagnetic correlations. $U=5$ in all panels.
Figure 5: Link correlations for $\epsilon=0.30$ along directions $\theta=\pi/3$ (left panels, $E_{0}$ maxima) and $\theta=\pi/6$ (right, $E_{0}$ minima). Upper and lower panel rows correspond to $U=10$ and $U=20$, respectively. Compared to Fig. 4(b) and (c), the correlations here are much more pronounced, with little difference between $U=10$ and 20.
@@
We now turn our attention to the effect of strain on the spin-spin correlations in the kagome ribbon. Figure 4 shows the heatmap of ${\cal C}_{\rm link}$ for angles where $E_{0}$ and $\zeta$ have maxima (i.e., higher frustration, $\theta=0$, $\pi/3$ and $32\pi/$ 3, left panels) or minima (lower frustration, $\theta=\pi/6$, $\pi/2$ and $65\pi/$ 6, right panels). We note that ${\cal C}_{\rm link}$ exhibits a structure of connected oblique lattices for strain directions where $\zeta$ is maximum (left panels). The links are homogeneous in strength/intensity throughout, with links stretched the most (those along $\theta$) having nearly vanishing correlation strength. In contrast, for angles where $E_{0}$ is minimum (right panels), the ${\cal C}_{\rm link}$ distribution appears as quasi-independent 1D chains of strong correlation links at angles orthogonal to the strain direction $\theta$.
@@
These features in the overall link distribution are even more pronounced for larger values of the Coulomb interaction $U$, as charge fluctuations are suppressed and antiferromagnetic correlations are enhanced overall. This is evident in Fig. 5 that shows link-correlations for $U=10$ (upper) and $U=20$ (lower panels). Left and right panels refer to $\theta=\pi/3$ and $\pi/6$, which compare with Fig. 4(c) and (b), respectively. Note that the Neel-like lines (right panels) and the oblique 2D lattice structures (left) are much more pronounced in this figure than in Fig. 4. We also note relatively little variation between $U=10$ and 20.
@@
While $\zeta(\theta)$ is a good indicator of the values of $\theta$ that most affect the ground state energy $E_{0}$, it does not directly quantify the spin frustration that may be present in the system. In an attempt to quantify the frustration content of the ground state, we introduce the quantity,
@@
which accounts for the spin-spin correlations within a given unit cell of the kagome system. In this expression, $J_{mn}=4t_{mn}^{2}/U$ represents the exchange coupling between neighbor spins $(m,n)$ within the unit cell considered, $\mathbb{C}_{B}^{j}$, and included in the sum. The unit cell used is shown by dashed magenta lines in Fig. 1. Finally, $\tilde{J}_{j}\equiv\sum_{(m,n)\in\mathbb{C_{B}^{\text{j}}}}J_{mn}$. In this definition, $f_{j}^{G}$ accounts for the couplings (including their the sign) between each pair of spins in all unit cells. Notice that $f_{j}^{G}\rightarrow 1$ for uncorrelated cells, while $f_{j}^{G}\rightarrow 0$ for both fully ferro- and antiferromagnetic correlated links.
@@
Figure 6 shows the average of this indicator over the entire system, $f^{G}(\theta)$, for three distinct values of $U$, over the range $20<\theta<\pi/$ 2; apart from small end-effect asymmetry, similar results are obtained for $\pi/2<\theta<\pi$. Notice that for large $U=10$ and 20, where antiferromagnetic correlations are better defined in the system, $f^{G}$ drops in value overall and exhibits well-pronounced maxima and minima that agree well with the corresponding behavior of $E_{0}(\theta)$ in Fig. 3. The nearly flat behavior of $f^{G}(\theta)$ for $U=5$ is likely associated with possible charge fluctuations in the system, suppressing local moments in lattice sites for weak interactions. This indicator is then not as effective in providing a quantitative assessment in the weak interaction regime, despite the clear impact of strain in the ${\cal C}_{\rm link}$ distribution.

### v2

@@ Abstract
We study the physics of the strong-coupling Hubbard model in a kagome lattice ribbon under mechanical tension and half-filling. It is known that in the absence of strain, the lattice symmetry of the system and strong electronic interactions induce magnetic frustration. As uniaxial strain is applied, the ribbon exhibits various configurations with energy oscillations that depend on the direction of the strain axis. The ground states are obtained by density-matrix renormalization-group calculations. We find that the system is characterized by strong antiferromagnetic bonds distributed throughout the lattice in directions and patterns that depend on strain directions and may coexist with easily polarizable sites that are only weakly correlated to their neighbors. We identify frustration and correlation measures that follow the strain and interaction dependence of the system well. These results illustrate that strain-dependent magnetic susceptibility could be explored experimentally to help probe the role of symmetry and interactions in these systems.
@@ I Introduction
The quantum spin liquid is an exotic and fascinating phase of matter [1, 2, 3]. Proposed by Anderson in [4], he suggested that such a phase could emerge as the ground state of a spin-1/2 antiferromagnet on geometrically frustrated triangular lattices [5]. However, after years of debate on whether a quantum spin-liquid (QSL) phase would emerge in a triangular antiferromagnet, detailed studies suggest this is not the case (see [1] and references therein). Indeed, the search for the realization of this elusive phase of matter has elicited great effort from the scientific community over the years [6, 7, 8, 9, 10].
@@
Different phases have been predicted theoretically in frustrated kagome lattices using antiferromagnetic Heisenberg and/or Hubbard models [40]. While the former accounts strictly for the spin degrees of freedom in the system, the latter also accounts for charge fluctuations and can then describe a wider range of phenomena. Schnyder et al. [17] showed that different ground states could be obtained by controlling the spatial anisotropy of antiferromagnetic Heisenberg couplings on a kagome lattice, resulting in spiral phases and possible noncoplanar order. More recently, Nayga and Vojta [41] have also shown that the ground state of a kagome magnet with classical spins can be controlled by applied mechanical distortion of the lattice, including the finding that triaxial strain results in a noncoplanar spin-liquid phase. These studies point out that the magnetic properties of kagome systems are very sensitive to changes in the geometric frustration of the lattice. The reduction in symmetry would then be expected to strongly affect systems that incorporate charge fluctuations, such as those present in kagome metals. The Hubbard model provides a natural description, as the relative hoppings can be seen as effective control parameters that allow one to address the important question of how strain modifies the ground state properties. Sun and Zhu [42] have shown recently that by controlling the Coulomb repulsion in a Hubbard model of a kagome ribbon it is possible to drive the system across multiple quantum phase transitions. This includes a strongly interacting regime where the ground state is effectively governed by an antiferromagnetic Heisenberg model and displays QSL characteristics. For weaker interaction, these authors saw evidence of different correlated phases in this geometrically frustrated system. We are interested in exploring the role that strain plays in the behavior of kagome systems as interactions change their response from the metallic to the insulating regime. The question we want to explore is how strain modifies possible frustrated ground states in the different regimes with varying correlations and charge fluctuations.
In this work, we investigate the effect of a uniaxial uniform strain in a kagome nanoribbon. A schematic representation of the system is depicted in Fig. 1. Adopting a single-band Hubbard model, the effective role of strain is to modify the hopping matrix elements between electronic orbitals localized at the lattice sites. This generally results in anisotropic couplings along the three lattice vectors, thus reducing geometric frustration in the system. Furthermore, varying the Coulomb repulsion allows glancing at how magnetic properties of the metallic and insulating ground states differ, thereby highlighting the relevance of charge fluctuations.
@@
By performing density-matrix renormalization-group (DMRG) calculations within a tensor networks platform [43, 44], we compute the ground state properties of the system, including the spatial distribution of spin-spin correlations and the local magnetization upon application of weak magnetic fields. Our results show that the anisotropy produced by the applied strain induces the formation of Neel lines of strongly antiferromagnetically correlated bonds arranged on different patterns and along different lattice directions, depending on the strain orientation. This behavior can be understood as arising from the suppression of the geometrical frustration in the system. Interestingly, we also find that scattered among the well-defined Neel lines with antiferromagnetic correlations, sites exist that are only weakly coupled to their neighbors. These interstitial sites are thus easily polarizable by weak external magnetic fields, while the Neel lines remain in their antiferromagnetic coupled structure. The pattern of Neel bonds and spatial distribution of polarizable sites depend critically on strain orientation and becomes more pronounced for larger values of Coulomb interaction as the system approaches the kagome antiferromagnet regime. To better characterize the response to strain, we introduce two intuitive quantities: a geometric indicator to quantify the structural frustration in the system and another that measures how local spin correlations reflect the degree of frustration in the ground state. These measures are good predictors of the different spin correlation profiles seen in the system as the strain and interaction change.
@@ II Model and numerical Methods
Figure 1: Schematic representation of the kagome nanoribbon corresponding to $L_{x}=4$ unit cells in the $x$ direction and $L_{y}$ = 3 unit cells in the $\hat{\bf e}_{2}$ direction. Circles represent the atomic sites of a tight-binding model, while solid lines correspond to nearest-neighbor hopping matrix elements. Red dashed lines at the top and bottom indicate periodic boundaries, in which case the ribbon folds into what will be referred to as the long cylinder geometry. Directions $\hat{\bf e}_{1}$ and $\hat{\bf e}_{2}$ span the unit cell, while the ${\boldsymbol{a}}_{j}$ ($j=1,2,3$) represent the nearest-neighbor vectors. The angle $\theta$ defines the direction of applied strain with respect to $\hat{\bf e}_{1}$. The dashed magenta hexagon contains a minimal unit cell and associated bonds used in the text.
@@
$H=\varepsilon_{0}\sum_{j,\sigma}n_{j\sigma}-\sum_{\begin{subarray}{c}<i,j>\ \sigma\end{subarray}}t_{ij}c_{i\sigma}^{\dagger}c_{j\sigma}+\sum_{j}\left(Un_{j\uparrow}n_{j\downarrow}+V_{z}S^{z}_{j}\right),$ (1)
@@ III Numerical Results
We set $t=1$ as the energy unit and $\hbar=1$ hereafter, and ensure that the system is in the half-filling regime by setting $\varepsilon_{0}=-U/2$. We explore values of $U$ from $U=5$ to $U=20$, corresponding to different phases of the system. For large values of $U$, the system is well described by an antiferromagnetic Heisenberg model [42]. The number of sites in the kagome nanoribbon increases rapidly with $L_{x}$ and $L_{y}$, $N=L_{y}(3L_{x}+2)$; we focus our analysis on a system with $L_{x}=12$ and $L_{y}=3$. Whenever the ribbon is folded along the $\boldsymbol{\hat{\text{e}}_{2}}$ direction into a cylinder, with sites periodically coupled as indicated by red dashed lines in Fig. 1, we call such arrangement the long cylinder geometry (LCG), which corresponds to the YC6 cylinder in Ref. [14]. On completely open boundaries, however, the configuration will be referred to as the flat ribbon geometry (FRG). We have allowed DMRG link dimension up to 5000, resulting in truncation errors typically less than 10-6.
@@ III.1 Unstrained vs strained regimes
We first look at the spin-spin correlations in the system for a strain-free FRG lattice, $\epsilon=0$. As defined above, we use ${\cal C}_{\rm link}$ to provide a convenient visualization of the correlations. Figure 2 shows a heat map of ${\cal C}_{\rm link}$ for different Hubbard interaction values, placing the system into different phases: $U=5$ (a), metallic phase; $U=10$ (b) and $U=20$ (c), quantum spin liquid/kagome antiferromagnetic phase [42]. As the color bar shows, darker links correspond to stronger antiferromagnetic pairs. For small $U$, the correlation links are not as strong and are homogeneously distributed, as the weak interactions in this metallic regime would suggest. Despite the homogeneous distribution of hopping terms in the Hamiltonian, we notice some differences, likely due to edge effects on the ribbon. For larger $U$, the antiferromagnetic correlations appear stronger (darker links) and not as homogeneously distributed as the system approaches the Heisenberg antiferromagnetic regime. In an infinite system, the spin-dominated ground state is expected to have high frustration without strain. A spin-liquid phase may even emerge, with not yet fully understood characteristics [14, 42]. We aim to study the role that strain may play in modifying the frustration otherwise present in the system. We will contrast Fig. 2 with the link correlation in the presence of strain to expose some of its effects on the ground state. All link correlations plots in this paper are for FRG.
@@
Figure 2: Link correlations, ${\cal C}_{\rm link}$, in the absence of strain ($\epsilon=0$) for (a) $U=5$, (b) $U=10$ and (c) $U=20$. Darker links represent stronger antiferromagnetic alignments, clearly enhanced in panel (c). The uniform link distribution in (a) becomes more inhomogeneous for increasing $U$. Apart from strong links along the edges of this flat ribbon, there appears to be no overall order of the antiferromagnetic bonds in (c).
@@
Figure 3: (a) Ground state energy per site, $E_{0}/N$, as function of strain direction $\theta$ for $\epsilon=0.1$ (a), $\epsilon=0.15$ (b), $\epsilon=0.2$ (c) and $\epsilon=0.3$ (d). Blue and red curves correspond to Flat Ribbon Geometry and Long Cylinder Geometry respectively. The black dashed line shows $\zeta(\theta)$ defined in Eq. (5) and scaled as per [46]. In all panels $U=5$, i.e., in the metallic regime for the unstrained system.
Figure 4: Link correlations for $\epsilon=0.30$ and different strain orientations: $\theta=0$ (a), $\theta=\pi/6$ (b), $\theta=\pi/3$ (c), $\theta=\pi/2$ (d), $\theta=2\pi/3$ (e) and $\theta=5\pi/6$ (f). The left (right) column panels correspond to maxima (minima) in $E_{0}(\theta)$ of Fig. 3. For energy maxima (left column), links with antiferromagnetic correlations form 2D oblique Lieb-like lattices. For energy minima (right column), links form 1D-like arrays with stronger (darker) antiferromagnetic correlations. $U=5$ in all panels. The top-right shows an inset with nearest-neighbor vectors to help identify the link orientations.
Figure 5: Link correlations for $\epsilon=0.30$ along $\boldsymbol{a_{1}}$ ($\theta=\pi/3$, left panels, a maximum in $E_{0}(\theta)$) and orthogonally to $\boldsymbol{a_{2}}$ ($\theta=\pi/6$, left panels, a minimum in $E_{0}(\theta)$). These configurations occur whenever strain is applied along $\boldsymbol{a_{1}}$ and orthogonally to ${\boldsymbol{a_{2}}}$, in this order (see top-right inset). Upper and lower panel rows correspond to $U=10$ and $U=20$, respectively. Compared to Fig. 4(b) and (c), the correlations here are much more pronounced, with little difference between $U=10$ and 20.
@@
We now analyze the effect of strain on the link-correlation pattern over the ribbon. Because the effect of strain on the correlations depends on the angle $\theta$, it is instructive to first analyze how the ground state energy $E_{0}$ changes with $\theta$. Figure 3 shows $E_{0}(\theta)$ vs $\theta$ for $U=5$ and increasing strain values. Blue circles and red squares correspond to FRG and LCG, respectively. In all cases, $E_{0}$ exhibits an oscillatory behavior with maxima and minima as $\theta$ varies from 00 to $\pi$, as well as an overall increasing value for larger strains. The oscillations with strain direction are better defined for $\epsilon=0.30$ [Fig. 3(d)] but already visible for $\epsilon=0.15$ [see Fig. 3(b)]. The minima occur for $\theta\simeq\pi/6$, $\pi/2$ and $65\pi/$ 6, which correspond to strain applied orthogonal to one of the lattice vectors ${\boldsymbol{a}}_{j}$. In contrast, the $E_{0}$ maxima are observed around $\theta=0$, $\pi/3$ and $32\pi/$ 3, for which the strain direction lies along one of the lattice vectors. The general behavior of $E_{0}$ increasing with $\epsilon$ can be seen to be associated with the suppression of antiferromagnetic correlations, which contribute to lowering the energy of the system. This is consistent with the fact that for LCG $E_{0}$ is lower, as additional singlet-like correlations appear across the edges. We also notice that $E_{0}(\theta)$ is not fully symmetric around $\theta=\pi/2$, a direct consequence of the slanted geometry that defines the ends of the ribbon. Such asymmetry is even more pronounced for LCG, as seen in the red curves of Fig. 3. This feature becomes clear once one realizes that, for LCG, configurations (e) and (f) in Fig. 4 loop ${\boldsymbol{a}}_{1}$ lines onto themselves (see Fig. 1), in great contrast to what happens with the remaining ones: for example, ${\boldsymbol{a}}_{2}$ lines are alternatingly coupled with one another in (b) and (c).
@@
This quantity can be scaled and compared with the ground state energy $E_{0}(\theta)$  [^1] The scaling factor is obtained as follows: with $\zeta_{max}\equiv\text{max}\left\{\zeta(\theta)\right\}$ and $\zeta_{min}\equiv\text{min}\left\{\zeta(\theta)\right\}$; the equivalent minimum (maximum) value of $E_{0}(\theta)$ for those angles in the FRG curve are $E_{min}$ ($E_{max}$). Then $\zeta\rightarrow\left(\frac{\zeta-\zeta_{min}}{\zeta_{max}-\zeta_{min}}\right)\left(E_{max}-E_{min}\right)+E_{min}$ gives the desired scale for $\zeta(\theta)$ shown in Fig. 3.. Figure 3 shows $\zeta(\theta)$ as a dashed black curve; evidently, the scaled function $\zeta(\theta)$ follows $E_{0}(\theta)$ quite well. As mentioned above, the results for LCG are more affected by end effects, and the agreement is poorer. Nevertheless, the positions of maxima and minima predicted by $\zeta(\theta)$ agree quite nicely with $E_{0}(\theta)$ for both FRG/LCG configurations obtained via DMRG. These results suggest that the maxima of $E_{0}(\theta)$ are obtained whenever the geometric frustration in the system is higher, even for the case $U=5$ (metallic regime when unstrained) shown in Fig. 3.
@@
We now turn our attention to the effect of strain on the spin-spin correlations in the kagome ribbon. Figure 4 shows the heatmap of ${\cal C}_{\rm link}$ for angles where $E_{0}$ and $\zeta$ have maxima (i.e., higher frustration, $\theta=0$, $\pi/3$ and $32\pi/$ 3, left panels) or minima (lower frustration, $\theta=\pi/6$, $\pi/2$ and $65\pi/$ 6, right panels). We note that ${\cal C}_{\rm link}$ exhibits a structure of connected oblique Lieb-like lattices [47] for strain directions where $\zeta$ is maximum (left panels). The links are homogeneous in strength/intensity throughout, with links stretched the most (those along $\theta$) having nearly vanishing correlation strength. In contrast, for angles where $E_{0}$ is minimum (right panels), the ${\cal C}_{\rm link}$ distribution appears as quasi-independent 1D chains of strong correlation links at angles orthogonal to strain direction $\theta$.
@@
These features in the overall link distribution are even more pronounced for larger values of the Coulomb interaction $U$, as charge fluctuations are suppressed and antiferromagnetic correlations enhanced overall. This is evident in Fig. 5 that shows link-correlations for $U=10$ (upper) and $U=20$ (lower panels). Left and right panels refer to $\theta=\pi/3$ and $\pi/6$, which compare with Fig. 4(c) and (b), respectively. Note that the Neel-like lines (right panels) and the oblique 2D lattice structures (left) are much more pronounced in this figure than in Fig. 4. We also note relatively little variation between $U=10$ and 20.
@@
While $\zeta(\theta)$ is a good indicator of $\theta$ values that most affect the ground state energy $E_{0}$, it does not directly quantify the spin frustration that may be present in the system. [Yet another indicator is discussed in Appendix B. A useful way to quantify the frustration content of the ground state is via the following quantity,
@@
which accounts for the spin-spin correlations within a given unit cell of the kagome system. In this expression, $J_{mn}=4t_{mn}^{2}/U$ represents the exchange coupling between neighbor spins $(m,n)$ within the unit cell considered, $\mathbb{C}_{B}^{j}$, and included in the sum. The unit cell used is shown by dashed magenta lines in Fig. 1. Finally, $\tilde{J}_{j}\equiv\sum_{(m,n)\in\mathbb{C_{B}^{\text{j}}}}\left|J_{mn}\right|$. In this definition, $f_{j}^{G}$ accounts for the couplings (including their the sign) between each pair of spins in all unit cells. Notice that $f_{j}^{G}\rightarrow 1$ for uncorrelated cells, while $f_{j}^{G}\rightarrow 0$ for both fully ferro- and antiferromagnetic correlated links.
Figure 6 shows the average of this indicator over the entire system, $f^{G}(\theta)$, for three distinct values of $U$, over the range $20<\theta<\pi/$ 2; apart from small end-effect asymmetry, similar results are obtained for $\pi/2<\theta<\pi$. Notice that for large $U=10$ and 20, where antiferromagnetic correlations are better defined in the system, $f^{G}$ drops in value overall and exhibits well-pronounced maxima and minima that agree well with the corresponding behavior of $E_{0}(\theta)$ in Fig. 3. The nearly flat behavior of $f^{G}(\theta)$ for $U=5$ is likely associated with charge fluctuations in the system, suppressing local moments in lattice sites for weak interactions. This indicator is then not as effective in providing a quantitative assessment in the weak interaction regime, despite the clear impact of strain on the ${\cal C}_{\rm link}$ distribution.
@@
Figure 7: Magnetization $M^{z}_{j}$ for $V_{z}=0.05$ and $\theta=0$ (a), $\theta={\pi}/{6}$ (b), $\theta={\pi}/{3}$ (c), $\theta={\pi}/{2}$ (d), $\theta={2\pi}/{3}$ (e) and $\theta={5\pi}/{6}$ (f). For all panels, we set $U=5$ and $\epsilon=0.30$. Faint dots on the right panels (where $E_{0}$ has minima in Fig. 3) feature sites belonging to Neel-like lines, while dark red dots correspond to loose sites readily polarized by the field. Top-right inset helps identifying, e.g., the red lines in (c) as being along the strain direction $\boldsymbol{a_{1}}$.
Figure 8: Local magnetization $M^{z}_{j}$ for $V_{z}=0.05$ and strain applied along $\theta=\pi/3$ (left) and $\theta=\pi/6$ (right). Upper and lower panels show results for $U=10$ and $U=20$, respectively. As in Fig. 7, faint color dots on the right panels represent sites belonging to Neel AF lines, while dark red dots correspond to loose sites more easily polarized by the field. Notice how the most energetically favorable top-left to bottom-right red-then-blue pattern results along $a_{2}$-connected spins. For all pannels, $\epsilon=0.30$.